OpenStudy (anonymous):

the gateway arch in st luis was built in 1965. it is the tallest manument in the united states. the arch can be modeled with the function y=-0.00635x^2+4x.

OpenStudy (anonymous):

how high above the ground is the tallest point of the arch and how far apart are the legs of the arch at their base

OpenStudy (anonymous):

h= -0.00635(x^2 - 630x) ... I rounded here max when x = 630/2 = 315 max height is -0.00635(315^2 - 630(315)) = approx 630m

OpenStudy (anonymous):

when width is 2 x 315 = 630 m

OpenStudy (anonymous):

or feet ??

OpenStudy (anonymous):

or set -.00635x^2 + 4x = 0 solves to give 630 ... the width find half of it 315 and subst in to get max height

OpenStudy (anonymous):

ok??